### v2

@@ 3.5 Phase3-4 : Hierarchical Tree Construction & Visualization
Step 1 - Root node determine:
Sort all the papers in descending order based on the quantity they have been cited. Select the top N papers with the highest citation counts as the root. This operation ensures that the root node has a high level of inheritability throughout the paper library.
@@
Step 1 - Root node determine:
Calculate the average of similarity score for each paper in the relevance matrix. Select the top N papers with the highest similarity score as the root nodes. This operation ensures that the root node has a high correlation index throughout the paper library.
@@
Step 3 - Parent node update: Set each leaf node as a new parent node and repeat Step 2. If a root node is not have relevance chain (unable to generate a leaf), the branch is terminated.
@@
Rule 1:
The selected nodes cannot be selected again.
@@
To provide researchers with a visual understanding(Fig.1-Fig.2) of the guidance on research directions provided by these selected papers. We use pyvis to visualize KG based on the configuration of Table 1. In addition to the co-occurring vocabulary of insight sentences between papers, we also extract keywords specific to each paper to capture its unique concept. When selecting keywords for each paper in the KG, we combine the ‘insight content’ from all papers in the subtree to create the overall text. The ’insight content’ of a specific article within the subtree is used as the partial text. Tf-idf[13] method used to extract partial text keywords from the global text composed of multiple paper-texts in a tree, and then display them as paper keywords in the node titles, and a part of the representative sample branch included insight sentences is shown in Fig.LABEL:figure3.
4 Conclusion
This study developed two type of hierarchical tree-structured knowledge graph called inheritance tree and relevance tree to support insight survey for research beginners. The process consists of four stages: data processing, insight sentence extraction, hierarchical tree construction, and KG visualization. First, we conducted high-dimensional secondary development of the S2ORC dataset to create an insight survey dataset that includes citation information and insight content. Then, using machine learning techniques, we extracted sentences from the insight survey dataset that express insight viewpoints on ’Issue finding’ and ’Issue resolved’. These sentences were parsed and used to construct a relevance matrix. Finally, based on the citation information and relevance matrix, we generated two types of hierarchical tree structures: Inheritance tree and Relevance tree. The generated KG demonstrate their rationality, indicating that they can provide key-information to assist researchers in gaining insights into the directions of the research topic. The knowledge graph also exhibits interpretability and potential for further development. For the future expansion and improvement of this study, the following points are proposed:
(1) Incorporate Additional Text for Text Similarity and relevance Chain Computation: Currently, this study only utilizes the text from the ‘conclusion’ and ‘limitation’ section. However, the Abstract and Introduction sections, specifically the part discussing previous issues, contain novel elements related to previous research topics and the subject of this study. Therefore, integrating these texts for high-dimensional training can optimize the content in knowledge graph.
(2) We conducted an objective evaluation of the classification accuracy of the ’Issue finding’ and ’Issue resolved’ viewpoints. However, integrating subjective evaluations from researchers into the generated knowledge graph remains a future challenge.

### v5

@@ III-D Phase3-4 : Hierarchical Tree Construction & Visualization
Step 1 - Root node determine: Sort all the papers in descending order based on the number of other papers that have cited them. Select the top N papers with the highest citation counts as the root. This operation ensures that the root node has a high level of inheritability throughout the paper library.
@@
Step 1 - Root node determine: Calculate the average similarity scores for each paper in the relevance matrix. Select the top N papers with the highest similarity score as the root nodes. This operation ensures that the root node has a high correlation index throughout the paper library.
@@
Step 3 - Parent node update: Set each leaf node as a new parent node and repeat Step 2. If a root node does not have a relevance chain (unable to generate a leaf), the branch is terminated.
@@
Rule 1: The selected nodes cannot be selected again.
@@
IV Conclusion
This study developed two types of hierarchical tree-structured knowledge graphs called inheritance tree and relevance tree to support insight surveys for research beginners. The process consists of four stages: data processing, insight sentence extraction, hierarchical tree construction, and KG visualization. First, we conducted high-dimensional secondary development of the S2orc dataset to create an insight survey dataset that includes citation information and insight content. Then, we extracted sentences from the insight survey dataset that express insight viewpoints on Issue finding’ and ’Issue resolved’ using machine learning techniques. These sentences were parsed and used to construct a relevance matrix. Finally, based on the citation information and relevance matrix, we generated two types of hierarchical tree structures: Inheritance and Relevance tree. The generated KG demonstrates their rationality, indicating that they can provide key-information to assist researchers in gaining insights into the directions of the research topic. The knowledge graph also exhibits interpretability and potential for further development. For the future expansion and improvement of this study, the following points are proposed:
(1) Incorporate Additional Text for Text Similarity and Relevance Chain Computation: Currently, this study only utilizes the content from the ‘conclusion’, ‘discussion’, and ‘limitation’ sections. However, the sections ‘abstract’ and ‘introduction,’ specifically the part discussing previous issues, contain insight elements related to ‘issue need to be solved in previous research’ and the ‘objective in this study’. Therefore, integrating these texts for high-dimensional training can optimize the content in the knowledge graph.
(2) We objectively evaluated the classification accuracy of the ’Issue Finding’ and ’Issue Resolved’ viewpoints. However, a future challenge is incorporating researchers’ subjective evaluations into the generated knowledge graph.